OpenStudy (anonymous):

Simplify: (sin Θ − cos Θ)2 + (sin Θ + cos Θ)2

OpenStudy (anonymous):

I'm one question away from finishing algebra 2 and I have no idea how to solve this...please help!

OpenStudy (anonymous):

sin and cos are just variables in this situation, you can think of them as y and x

OpenStudy (anonymous):

so 4 sin \[\theta\]

OpenStudy (solomonzelman):

\(\large\color{blue}{(\sin^2x −2\sin x \cos x+ \cos^2x) + (\sin^2x +2\sin x \cos x+ \cos^2x) }\) \(\large\color{blue}{\sin^2x −2\sin x \cos x+ \cos^2x + \sin^2x +2\sin x \cos x+ \cos^2x }\) \(\large\color{blue}{2\sin^2x + 2\cos^2x }\) \(\large\color{blue}{2(\sin^2x + \cos^2x) }\) \(\large\color{blue}{2(1) }\) \(\large\color{blue}{2. }\)

OpenStudy (solomonzelman):

question marks, god dang it!!

OpenStudy (mathstudent55):

Use the square of binomial patterns below to square each binomial. Then combine like terms. \((a + b)^2 = a^2 + 2ab +b^2\) \((a - b)^2 = a^2 - 2ab +b^2\)

OpenStudy (anonymous):

@SolomonZelman is that all is needed to solve it? If it is, thank you SO much!!

OpenStudy (solomonzelman):

ye,s expand, add like terms, and use the Pythagorean identity.